# On an Optimal Stopping Problem with a Discontinuous Reward

### v2

@@ 4.4. Characterization of the Surrender Region
The next corollary shows that when ${L(t,x)\geq 0}$ for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$, then $T$, which was shown to be an optimal stopping time for (7) in Proposition 3.1, is indeed unique.
Corollary 4.3.
If $L(t,x)\geq 0$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, then $\mathcal{S}=\emptyset$ and $T$ is the unique optimal stopping time for (7).
Proof.
From Proposition 3.1, we have that $v(t,x)=h(t,x)$, for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$. Now using Itô’s lemma on the discounted surrender value process and the zero-mean property of the stochastic integral, we find that for any $(t,x)\in[0,T)\times\mathbb{R}_{+}$,
$v(t,x)$ $=\mathbb{E}\left[e^{-r(T-t)}\max(G,F_{T}^{t,x})\right]$
$>g(t,x)x.$
The shape of the surrender region can vary significantly; see for example Figures 4 and 5 of MacKay et al. (2017). However, when the fee rate is constant and the (time-dependent) surrender charge function satisfies certain conditions, MacKay (2014), Appendix 2.A shows that the surrender region is a connected set. Proposition 4.8 below shows that as soon as the surrender region is non-empty and the fee and surrender charge functions depend only on time, then $t$-section of $\mathcal{S}$ has the form $\mathcal{S}_{t}=[b(t),\infty)$ for some $b(t)\in\mathbb{R}_{+}\cup\{\infty\}$.
Proposition 4.8 builds on Jacka (1991), Proposition 2.1 in the context of a bounded and time-homogeneous reward function in a time-homogeneous market model.
Proposition 4.8.
Let Assumptions 4.1, 4.2, and 4.3 hold. If $L(t)<0$ for all $t\in[0,T)$ then the $t$-sections of $\mathcal{S}$, denoted by $\mathcal{S}_{t}$, are of the form
Proof.
Let $L(t)<0$ for all $t\in[0,T)$. It follows from Proposition 4.6 (ii) that $\mathcal{S}\neq\emptyset$, and thus there exists $t\in[0,T)$ for which $\mathcal{S}_{t}\neq\emptyset$. Fix $t\in[0,T)$. If $\mathcal{S}_{t}=\emptyset$ then $b(t)=\infty$ and the proof is complete. If $\mathcal{S}_{t}\neq\emptyset$, define
@@
since $g$ takes values in $(0,1]$. Next, we show that $\mathcal{S}_{t}=[b(t),\,\infty)$. Fix $t\in[0,\,T)$ and note that if $\mathcal{S}_{t}=[b(t),\infty)$, then $\mathcal{C}_{t}=(0,b(t))$. Thus, we need to prove that $x\in\mathcal{C}_{t}\Rightarrow y\in\mathcal{C}_{t}$ for any $y<x$, which we do next. Recall that $\tau_{t}^{x}=\inf\left\{t\leq s\leq T\,\big{|}\,v(s,F_{s}^{t,x})=\varphi(s,F_{s}^{t,x})\right\}$ is optimal for $v(t,x)$. Hence, we have that
@@
since $g(T)=1$. Combining (50) and (51) yields
@@
Finally, since $x\in\mathcal{C}_{t}$, we have $v(t,\,x)>g(t)x$, so that $v(t,y)>g(t)y$, which implies $y\in\mathcal{C}_{t}$. ∎
Under the assumptions of Proposition 4.8, $b(t)$ is the smallest sub-account value for which it is optimal to surrender the contract at time $t<T$, and for any fund value greater than $b(t)$, it is also optimal to surrender, so that
and
@@ 4.5. Equivalence of the Optimal Stopping Problems
Suppose Assumption 4.2 holds. If $L(t)<0$ for all $t\in[0,T)$, then $\mathcal{S}=\widetilde{\mathcal{S}}$.
@@
Under the assumptions of Proposition 4.9, ${v(t,x)>h(t,x)}$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, where ${h(t,x)=\mathbb{E}\left[e^{-r(T-t)}\max(G,F_{T}^{t,x})\right]}$.
@@
Using arguments similar to those of the proof of Lemma 4.4, we find a contradiction and therefore conclude that $v(t,x)>h(t,x)$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$. ∎
@@
To show that $\mathcal{S}\subseteq\widetilde{\mathcal{S}}$, fix ${(t,x)\in\mathcal{S}}$. Then by Lemma 3.6, ${\varphi(t,x)\geq h(t,x)}$, so that ${\varphi(t,x)=\max(\varphi(t,x),h(t,x))}$. By definition of $\mathcal{S}$, $v(t,x)=\varphi(t,x)=\max(\varphi(t,x),h(t,x))$, and thus $(t,x)\in\widetilde{\mathcal{S}}$. To show that $\mathcal{S}\supseteq\widetilde{\mathcal{S}}$, we show $\mathcal{C}\subseteq\widetilde{\mathcal{C}}$. Fix $(t,x)\in\mathcal{C}$, so that ${v(t,x)>\varphi(t,x)}$. Since $L(t)<0$ for all $t\in[0,T)$, we have by Lemma 4.6 that ${v(t,x)>h(t,x)}$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$. Thus, $v(t,x)>\max(\varphi(t,x),h(t,x))$, which confirms that $(t,x)\in\widetilde{\mathcal{C}}$. ∎
Combining the results of Lemma 3.5 and Proposition 4.9 yields the next corollary.
@@
Suppose Assumptions 4.1 (i) and 4.2 hold. If $L(t)<0$ for all $t\in[0,T)$, then $\tau_{t}^{x}=\widetilde{\tau}_{t}^{x}$.
@@
The proof follows immediately from Lemma 3.5, Proposition 4.9 and continuity of the value function in Theorem 4.1. ∎
Remark 4.15.
From the results above, we can conclude that when the fee and surrender charge functions only depend on time (Assumption 4.2) and satisfy other regularity assumptions (Assumptions 4.1 (i)), if $L(t)<0$ for all $t\in[0,T)$, the optimal stopping problem with the continuous reward function (14) is equivalent to the optimal stopping problem with the discontinuous reward function (7). That is, the two problems lead to the same value function (Theorem 3.4), surrender region (Proposition 4.9), and optimal stopping time (Corollary 4.4). It also follows that the optimal surrender boundary and the continuation region will be the same for the two problems. Numerical examples in Section 5 show that when these conditions are not satisfied the two problems may still lead to the same surrender region. This suggests that the conditions stated in Proposition 4.9 can be relaxed.
@@ 5. Numerical Examples
In this section, we give two simple examples to illustrate the differences between the optimal stopping problem involved in variable annuities pricing and the one stemming from American call and put option pricing often studied in the literature. When valuing variable annuities, the fee and the surrender charge functions can lead to a disconnected surrender region and a discontinuous optimal stopping boundary. We consider the time-dependent fee functions $c_{1}$ and $c_{2}$ defined by
$c_{1}(t)$ $=0.010908\times\mathbf{1}_{\{t\leq 5\}}+0.005454\times\mathbf{1}_{\{5<t\leq 10\}}+0.010908\times\mathbf{1}_{\{10<t\leq T\}},$
$c_{2}(t)$ $=0.010908\times\mathbf{1}_{\{t\leq 10\}}+0.005454\times\mathbf{1}_{\{10<t\leq T\}},$
with $T=15$. The surrender function is set to $g(t)=e^{-\kappa(T-t)}$, with $\kappa=0.0055$. A simple calculation shows that $L(t)>0$ when $t\in(5,10]$ for $c_{1}$ and $t\in(10,T]$ for $c_{2}$.
@@
(a) Discontinuous reward function with $c_{1}$ (b) Continuous reward function with $c_{1}$
Refer to caption Refer to caption
(c) Discontinuous reward function with $c_{2}$ (d) Continuous reward function with $c_{2}$
Figure 1. Comparison of the continuation regions (in red) of the optimal stopping problems with the discontinuous and the continuous reward functions (7) and (14), respectively. The value of the variable annuity contract is approximated using the continuous-time Markov chain approximation described in MacKay et al. (2023). Market and VA parameters are $r=0.03$, $\sigma=0.2$, $F_{0}=G=100$ and $T=15$.
Figure 1 compares the continuation regions (in red) of the two optimal stopping problems, with the discontinuous and the continuous reward functions, in (7) (left panel) and (14) (right panel). We first observe that the surrender regions (in white) are empty between year $5$ and $10$ for the fee function $c_{1}$, and between year $10$ and $15$ for $c_{2}$. This corresponds to the time intervals during which $L(t)>0$, as shown theoretically in Proposition 4.6 (i). Thus, the optimal surrender boundaries of the two problems (left and right panels) are discontinuous when the fee is given by $c_{1}$. It is already known from the literature of variable annuity that the surrender region can have different shapes depending on the fee function, see for example MacKay et al. (2017), Figures 4 and 5. However, such disconnected sets for the surrender region have not been observed in prior numerical work on similar problems, see among others Bernard et al. (2014b), MacKay et al. (2017), Kang and Ziveyi (2018), and MacKay et al. (2023). This simple example illustrates well how different the optimal stopping problem involved in variable annuities pricing differs from the traditional American call and put options often studied in the literature.
Another important feature that is highlighted in the second example (when the fee is modeled by $c_{2}$) is that we no longer have $\lim_{t\rightarrow T}b(t)=G$. This is because the surrender region is empty close to maturity (since $L(t)>0$).
Finally, we observe the two problems, with the discontinuous (left panel) and the continuous (right panel) reward functions, lead to the same continuation and surrender regions, despite the fact that $L(t)<0$ is not satisfied for all $t\in[0,T)$. This indicates that the conditions stated in Proposition 4.9, for the two problems to have the same surrender region, are only sufficient and can be further relaxed. Finding the necessary conditions that make the two optimal stopping problems equivalent is left as future research.
@@ 6. Conclusion
In particular, the continuous reward representation allows us to use existing results from American option pricing theory to obtain the continuity of the value function. We also prove its convexity in the state variable, and its Lipschitz property in $x$ and (locally) in $t$. Then, we derive an integral expression for the early surrender premium, generalizing the results of Bernard et al. (2014a) to general time-dependent fee and surrender charge functions. This second representation of the value function is also known as the early exercise premium representation in the American option pricing literature. From there, we develop a third representation for the value function in terms of the current surrender value and an integral expression that only takes value in the continuation region. We call this third representation the continuation premium representation.

### v3

@@ 4.4. Characterization of the Surrender Region
Propositions 4.6 and 4.7 can now be used to prove Theorem 4.3.
Theorem 4.3.
To show part (i), let $L(t)<0$ for all $t\in[0,T)$. It follows from Proposition 4.6 (ii) that $\mathcal{S}\neq\emptyset$, and thus there exists $t\in[0,T)$ for which $\mathcal{S}_{t}\neq\emptyset$. Fix $t\in[0,T)$. If $\mathcal{S}_{t}=\emptyset$ then $b(t)=\infty$ and the proof is complete. If $\mathcal{S}_{t}\neq\emptyset$, define
@@
since $g$ takes values in $(0,1]$.
Next, we show that $\mathcal{S}_{t}=[b(t),\,\infty)$. Fix $t\in[0,\,T)$ and note that if $\mathcal{S}_{t}=[b(t),\infty)$, then $\mathcal{C}_{t}=(0,b(t))$. Thus, we need to prove that $x\in\mathcal{C}_{t}\Rightarrow y\in\mathcal{C}_{t}$ for any $y<x$, which we do next.
Recall that $\tau_{t}^{x}=\inf\left\{t\leq s\leq T\,\big{|}\,v(s,F_{s}^{t,x})=\varphi(s,F_{s}^{t,x})\right\}$ is optimal for $v(t,x)$. Hence, we have that
@@
since $g(T)=1$. Combining (LABEL:eqMaxF1) and (49) yields
@@
Finally, since $x\in\mathcal{C}_{t}$, we have $v(t,\,x)>g(t)x$, so that $v(t,y)>g(t)y$, which implies $y\in\mathcal{C}_{t}$.
For part (ii), we have from Proposition 3.1 that $v(t,x)=h(t,x)$, for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$. Now using Itô’s lemma on the discounted surrender value process and the zero-mean property of the stochastic integral, we find that for any $(t,x)\in[0,T)\times\mathbb{R}_{+}$,
$v(t,x)$ $=\mathbb{E}\left[e^{-r(T-t)}\max(G,F_{T}^{t,x})\right]$
$>g(t,x)x.$
@@ 4.5. Equivalence of the Optimal Stopping Problems
Let $L$ be defined by (18). If $L(t)<0$ for all $t\in[0,T)$, then $\mathcal{S}=\widetilde{\mathcal{S}}$.
@@
If $L(t)<0$, ${v(t,x)>h(t,x)}$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, where ${h(t,x)=\mathbb{E}\left[e^{-r(T-t)}\max(G,F_{T}^{t,x})\right]}$.
@@
Using arguments similar to those of the proof of Proposition 4.6(ii), we find a contradiction and therefore conclude that $v(t,x)>h(t,x)$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$. ∎
@@
To show that $\mathcal{S}\subseteq\widetilde{\mathcal{S}}$, fix ${(t,x)\in\mathcal{S}}$, and observe that ${\varphi(t,x)=v(t,x)\geq h(t,x)}$. Hence, ${v(t,x)=\max(\varphi(t,x),h(t,x))}$, and thus $(t,x)\in\widetilde{\mathcal{S}}$, by definition of $\widetilde{\mathcal{S}}$
To show that $\mathcal{S}\supseteq\widetilde{\mathcal{S}}$, we show $\mathcal{C}\subseteq\widetilde{\mathcal{C}}$. Fix $(t,x)\in\mathcal{C}$, so that ${v(t,x)>\varphi(t,x)}$. Since $L(t)<0$ for all $t\in[0,T)$, we have by Lemma 4.5 that ${v(t,x)>h(t,x)}$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$. Thus, $v(t,x)>\max(\varphi(t,x),h(t,x))$, which confirms that $(t,x)\in\widetilde{\mathcal{C}}$. ∎
The next corollary establishes the equality of $\tau_{t}^{x}$ and $\widetilde{\tau}_{t}^{x}$ under a simple condition, complementing the previous results of Lemma 3.3.
@@
If $L(t)<0$ for all $t\in[0,T)$, then $\tau_{t}^{x}=\widetilde{\tau}_{t}^{x}$.
@@
The stopping times $\tau^{x}_{t}$ and $\widetilde{\tau}^{x}_{t}$ can be written as
$\begin{array}[]{cc}\tau^{x}_{t}&=\inf\{t\leq s\leq T|(s,F_{s}^{t,x})\in\mathcal{S}\},\\ \widetilde{\tau}^{x}_{t}&=\inf\{t\leq s\leq T|(s,F_{s}^{t,x})\in\widetilde{\mathcal{S}}\}.\end{array}$
The proof then follows immediately from Proposition 4.8. ∎
From the results above, we can conclude that when the fee and surrender charge functions only depend on time (Assumption 4.2) and satisfy other regularity assumptions (Assumption 4.1), if $L(t)<0$ for all $t\in[0,T)$, the optimal stopping problem with the continuous reward function (13) is equivalent to the optimal stopping problem with the discontinuous reward function (7). That is, the two problems lead to the same value function (Theorem 3.4), surrender region (Proposition 4.8), and optimal stopping time (Corollary 4.3). It also follows that the optimal surrender boundary and the continuation region will be the same for the two problems.
To provide some intuition for this result, we remark that the condition $L(t)<0$ is associated with a non-empty surrender region (see Proposition 4.6); at any point during the life of the contract, there is always a possibility that early surrender will become optimal in the future. Th, enthe value of the VA contract is strictly above that of the EPV of the maturity benefit, and if surrender occurs it will be due to the contract value being equal to the surrender benefit. When this is the case, all three stopping times coincide (see the discussion after the proof of Lemma 3.3).
@@ 5. Numerical Examples
In this section, we present two simple examples to highlight the specific features of the optimal stopping problem stemming from variable annuity pricing, when compared to those associated with the American call and put option pricing. In particular, with variable annuities, the surrender region can be disconnected and the optimal stopping boundary, discontinuous.
We consider the time-dependent fee functions $c_{1}$ and $c_{2}$, given by
$c_{1}(t)$ $=\frac{1}{100}\left(0.00889t^{2}-0.1330t+1\right),$
$c_{2}(t)$ $=\frac{1}{100}\left(-0.0333t+1\right),$
for $0\leq t\leq T$, with $T=15$. The surrender charge function is set to $g(t)=e^{-\kappa(T-t)}$, with $\kappa=0.0055$, which is a commonly used form in the variable annuity literature. The resulting surrender charge at inception of the contract is close to $8\%$, and it goes down to less than $0.5\%$ at $t=14$, one year before maturity.
The fee functions were chosen to illustrate some of the results obtained in Section 4. Function $c_{1}$ shows the impact of a change in the sign of $L(t)$; it is negative on $[0,5.17]\cup[9.79,15]$ and positive elsewhere. Function $c_{2}$ was chosen such that $L(t)>0$ for $t\in(13.5,T]$ for $c_{2}$, thus eliminating the risk-neutral surrender incentive in the last 1.5 year of the contract. In both cases, the annual fee rates vary between $0.5\%$ and $1.0\%$. In practice, the fee is typically set as a constant and the surrender charge only depends on time. Numerical examples in this setting are provided, for example, in Bernard et al. (2014b), MacKay (2014) and MacKay et al. (2017).
@@
(a) $c_{1}$ (b) $c_{2}$
Figure 1. The continuation region (in red) of the optimal stopping problem with the discontinuous reward function (7). The value of the variable annuity contract is approximated using the continuous-time Markov chain approximation described in MacKay et al. (2023). Market and VA parameters are $r=0.03$, $\sigma=0.2$, $F_{0}=G=100$ and $T=15$.
Figure 1 shows the continuation regions (in red) associated to the optimal stopping problem in (7) for the two fee functions $c_{1}$ and $c_{2}$. The surrender regions (in white) are empty between years $5$ and $10$ for the fee function $c_{1}$, and between years $13.5$ and $15$ for $c_{2}$. This corresponds to the time intervals during which $L(t)>0$, illustrating a result of Proposition 4.7 (ii). Thus, the optimal surrender boundary is discontinuous when the fee is given by $c_{1}$. It is already known from the literature of variable annuity that the surrender region can have different shapes depending on the fee function, see for example MacKay et al. (2017), Figures 4 and 5. However, such disconnected sets for the surrender region have not been observed in prior numerical work on similar problems, see among others Bernard et al. (2014b), MacKay et al. (2017), Kang and Ziveyi (2018), and MacKay et al. (2023). This simple example illustrates important differences between the optimal stopping problem involved in variable annuities pricing and the one associated with American call and put options.
Another important feature highlighted in the second example (with fee function $c_{2}$) is that we no longer have $\lim_{t\rightarrow T}b(t)=G$. This is because the surrender region is empty for $t>13.5$, just before maturity (since $L(t)>0$). Moreover, in this particular case, since $\mathcal{S}\cap([13.5,T)\times\mathbb{R}_{+})=\emptyset$, we have $v(t,x)=h(t,x)$ for $(t,x)\in[13.5,T)$, so that $\widetilde{\mathcal{S}}\cap([13.5,T)\times\mathbb{R}_{+})=[13.5,T)\times\mathbb{R}_{+}$, providing another example in which the optimal stopping problems with the discontinuous and the continuous reward functions have different surrender (and continuation) regions.
Finally, we note that in our figures, the surrender region becomes empty slightly before $t=5$ for the fee function $c_{1}$, and before $t=13.5$ for $c_{2}$. This is likely due to the $y$-axis stopping at 500; it may be that there exists a surrender region at $t=5$ for account values that are (much) higher than 500, and that the surrender boundary is continuous in a general sense, “reaching infinity” continuously when the sign of $L$ changes. This would also mean that theoretically, we could show that $L(t)\leq 0$ if and only if $\mathcal{S}_{t}$ is non-empty. Verifying such properties of the surrender boundary is left for future research.
An economic interpretation for the condition $L(t)>0$, which leads to the surrender region being empty, is that the surrender charge to be paid upon surrender is greater than the expected future fee payment. Surrenders then become more expensive and the policyholder prefers to hold on to her contract. This is also discussed in MacKay et al. (2017) in a less general setting.
@@ 6. Conclusion
In particular, the continuous reward representation allows us to use existing results from American option pricing theory to obtain the continuity of the value function. We also prove its convexity in the state variable, its Lipschitz property in $x$ and characterize its continuity in $t$. Then, we derive an integral expression for the early surrender premium, generalizing the results of Bernard et al. (2014a) to general time-dependent fee and surrender charge functions. This second representation of the value function is also known as the early exercise premium representation in the American option pricing literature. From there, we develop a third representation for the value function in terms of the current surrender value and an integral expression that only takes value in the continuation region. We call this third representation the continuation premium representation.